# Client-Aided Secure Two-Party Computation of Dynamic Controllers

### v1

@@ 8 Conclusion
We proposed a two-party protocol for the computation of a dynamic controller. The proposed protocol was realized based on a secret sharing scheme without the need for controller state decryption or input re-encryption and achieved statistical security. The performance of the dynamic controller under the proposed protocol can be made arbitrarily close to that of the original controller by approximating the controller parameters and initial state as sufficiently precise fixed-point numbers. We compared the computational and communication complexities of the protocol with those of conventional encrypted controllers and presented some improvements to the protocol using coordinate transformation and cryptographic primitives. Furthermore, the feasibility of the protocol was demonstrated through numerical examples. In future work, we will extend the protocol to more advanced controls, such as model predictive control.
@@
where |$\absolutevalue*{\lceil m/2^{\ell}\rfloor+w}\leq 2^{\kappa-\ell-1}+1<q/2$. Therefore, the claim holds from (P1).
@@
$=\left\lceil\quantity(\tilde{A}\tilde{x}(t)+\tilde{B}\tilde{y}(t))/2^{\ell}\right\rfloor+w(t),$
$=A\tilde{x}(t)+B\bar{y}(t)+e(t)+w(t),$
where $w(t)\in\{-1,0,1\}^{n}$, and $e(t)=\lceil A\tilde{x}(t)+B\tilde{y}(t)\rfloor-(A\tilde{x}(t)+B\bar{y}(t))\in(-1/2,1/2]^{n}$. Let $\delta(t)=e(t)+w(t)$. Then, we obtain $\tilde{x}(t+1)=A\tilde{x}(t)+B\bar{y}(t)+\delta(t)$ with $\norm{\delta(t)}_{\infty}\leq 3/2$. Similary, if $\tilde{C}\tilde{x}(t)+\tilde{D}\tilde{y}(t)\in\mathbb{Z}_{q}^{m}$ and $\bar{y}(t)\in\mathbb{Z}_{q}^{p}$, it follows that
@@
$=\tilde{C}\tilde{x}(t)+\tilde{D}\tilde{y}(t)=\bar{C}\tilde{x}(t)+\bar{D}\bar{y}(t).$
@@
where $d_{1}(t)=2^{-\ell}\bar{y}(t)-\hat{y}(t)$ and $d_{2}(t)=2^{-\ell}\delta(t)$. Hence, we have

### v2

@@ 8 Conclusion
We proposed a two-party protocol for the computation of a dynamic controller. The proposed protocol was realized based on a secret sharing scheme without the need for controller state decryption or input re-encryption and achieved statistical security. The performance of the dynamic controller under the proposed protocol can be made arbitrarily close to that of the original controller by approximating the controller parameters and initial state as sufficiently precise fixed-point numbers. We compared the computational and communication complexities of the protocol with those of conventional encrypted controllers and presented some improvements to the protocol using coordinate transformation and cryptographic primitives. Furthermore, the feasibility of the protocol was demonstrated through numerical examples.
In future work, we will extend the protocol to more advanced control strategies, such as model predictive control. We will also consider guaranteeing security not only against semi-honest adversaries but also against active adversaries who deviate from our intended behavior. This would address the limitation imposed by Assumption 3.
@@
[…]
@@
$=\left\lceil\quantity(\bar{A}\tilde{x}(t)+\bar{B}\bar{y}(t))/2^{\ell}\right\rfloor+w(t).$
Similary, if $\tilde{C}\tilde{x}(t)+\tilde{D}\tilde{y}(t)\in\mathbb{Z}_{q}^{m}$ and $\bar{y}(t)\in\mathbb{Z}_{q}^{p}$, it follows that
@@
$=\bar{C}\tilde{x}(t)+\bar{D}\bar{y}(t).$
@@
where $d_{1}(t)=2^{-\ell}\bar{y}(t)-\hat{y}(t)$, $d_{2}(t)=2^{-\ell}(e(t)+w(t))$, and $e(t)=\lceil A\tilde{x}(t)+B\bar{y}(t)\rfloor-(A\tilde{x}(t)+B\bar{y}(t))$. Hence, we have